# High-dimensional Gaussian and bootstrap approximations for robust means

### v3

@@ Appendix B Preparatory lemmas
Let $\overline{\Sigma}_{\varepsilon_{n}}$ and $\underline{\Sigma}_{\varepsilon_{n}}$ be the matrices with $(j,k)$th entry
@@
Fix $c\in(1,\sqrt{1.5})$, and let Assumption 2.1 be satisfied with $m>2$. If $\varepsilon_{n}\in(0,0.5)$, with $\varepsilon_{n}$ as in (5), it holds that
$\max_{j,k\in\mathinner{\{1,\ldots,d\}}}\mathinner{\Bigl(\mathinner{\!\bigl\lvert\overline{\Sigma}_{\varepsilon_{n},j,k}-\Sigma_{j,k}\bigr\rvert}\vee\mathinner{\!\bigl\lvert\underline{\Sigma}_{\varepsilon_{n},j,k}-\Sigma_{j,k}\bigr\rvert}\Bigr)}\leq 32\sigma_{m}^{2}\mathinner{\Bigl(\frac{3c-1}{c-1}\Bigr)}\mathinner{\Bigl(\frac{c+1}{c}\Bigr)}\varepsilon_{n}^{1-\frac{2}{m}}.$ (B.1)
@@
$=T_{1,j}+T_{2,j}+T_{3,j}.$
Thus, for $j,k\in\mathinner{\{1,\ldots,d\}}$, it follows by the Cauchy-Schwarz inequality that
@@
Recall that $\overline{\alpha}_{j}=Q_{\varepsilon_{n}+c^{-1}\varepsilon_{n}}(X_{1,j})$ such that $\mathbb{P}(X_{1,j}<\overline{\alpha}_{j})\leq\varepsilon_{n}+c^{-1}\varepsilon_{n}$. Thus, it follows by Lemma G.1, followed by Hölder’s inequality, that
$\mathbb{E}T_{2,j}^{2}$ $\leq\mathbb{E}(\mu_{j}+\sigma_{m}/(1-\varepsilon_{n}-c^{-1}\varepsilon_{n})^{1/m}-X_{1,j})^{2}\mathds{1}\mathinner{\bigl(X_{1,j}<\overline{\alpha}_{j}\bigr)}$
$\leq\mathinner{\Bigl(\mathbb{E}\mathinner{\!\bigl\lvert\mu_{j}-X_{1,j}+\sigma_{m}/(1-\varepsilon_{n}-c^{-1}\varepsilon_{n})^{1/m}\bigr\rvert}^{m}\Bigr)}^{2/m}\mathbb{P}(X_{1,j}<\overline{\alpha}_{j})^{1-\frac{2}{m}}$
$\leq\mathinner{\Bigl(2^{m}\sigma_{m}^{m}+2^{m}\sigma_{m}^{m}/(1-\varepsilon_{n}-c^{-1}\varepsilon_{n})\Bigr)}^{2/m}(\varepsilon_{n}+c^{-1}\varepsilon_{n})^{1-\frac{2}{m}}$
$\leq 4\sigma_{m}^{2}\mathinner{\bigl(1+1/(1-\varepsilon_{n}-c^{-1}\varepsilon_{n})^{2/m}\bigr)}\mathinner{\Bigl(\frac{c+1}{c}\Bigr)}^{1-\frac{2}{m}}\varepsilon_{n}^{1-\frac{2}{m}}.$ (B.3)
By the same arguments, using $\mathbb{P}(X_{1,j}>\overline{\beta}_{j})=1-\mathbb{P}(X_{1,j}\leq\overline{\beta}_{j})\leq\varepsilon_{n}-c^{-1}\varepsilon_{n}$,
$\mathbb{E}T_{3,j}^{2}\leq 4\sigma_{m}^{2}\mathinner{\bigl(1+1/(1-\varepsilon_{n}+c^{-1}\varepsilon_{n})^{2/m}\bigr)}\mathinner{\Bigl(\frac{c-1}{c}\Bigr)}^{1-\frac{2}{m}}\varepsilon_{n}^{1-\frac{2}{m}}.$ (B.4)
@@
Recalling that $\overline{\alpha}_{j}=Q_{\varepsilon_{n}+c^{-1}\varepsilon_{n}}(X_{1,j})$, it follows by Lemma G.1, followed by Hölder’s inequality, that
@@
$\leq\mathbb{E}\mathinner{\bigl(|X_{1,j}-\mu_{j}||\mu_{k}+\sigma_{m}/(1-\varepsilon_{n}-c^{-1}\varepsilon_{n})^{1/m}-X_{1,k}|\mathds{1}\mathinner{\bigl(X_{1,k}<\overline{\alpha}_{k}\bigr)}\bigr)}$
$\leq\mathbb{E}\mathinner{\bigl(|X_{1,j}-\mu_{j}||X_{1,k}-\mu_{k}|+|X_{1,j}-\mu_{j}|\sigma_{m}/(1-\varepsilon_{n}-c^{-1}\varepsilon_{n})^{1/m}\bigr)}\mathds{1}\mathinner{\bigl(X_{1,k}<\overline{\alpha}_{k}\bigr)}$
$\leq\mathinner{\Bigl(\mathbb{E}\mathinner{\bigl(|X_{1,j}-\mu_{j}||X_{1,k}-\mu_{k}|+|X_{1,j}-\mu_{j}|\sigma_{m}/(1-\varepsilon_{n}-c^{-1}\varepsilon_{n})^{1/m}\bigr)}^{\frac{m}{2}}\Bigr)}^{\frac{2}{m}}(\varepsilon_{n}+c^{-1}\varepsilon_{n})^{1-\frac{2}{m}}$
$\leq\mathinner{\Bigl(2^{m/2}(\sigma_{m}^{m}+\sigma_{m/2}^{m/2}\sigma_{m}^{m/2}/(1-\varepsilon_{n}-c^{-1}\varepsilon_{n})^{1/2}\Bigr)}^{2/m}\mathinner{\Bigl(\frac{c+1}{c}\Bigr)}^{1-\frac{2}{m}}\varepsilon_{n}^{1-\frac{2}{m}}$
$\leq 2\sigma_{m}^{2}\mathinner{\bigl(1+1/(1-\varepsilon_{n}-c^{-1}\varepsilon_{n})^{1/m}\bigr)}\mathinner{\Bigl(\frac{c+1}{c}\Bigr)}^{1-\frac{2}{m}}\varepsilon_{n}^{1-\frac{2}{m}}.$ (B.5)
@@
$\mathinner{\!\bigl\lvert\mathbb{E}T_{1,j}T_{3,k}\bigr\rvert}\leq 2\sigma_{m}^{2}\mathinner{\bigl(1+1/(1-\varepsilon_{n}+c^{-1}\varepsilon_{n})^{1/m}\bigr)}\mathinner{\Bigl(\frac{c-1}{c}\Bigr)}^{1-\frac{2}{m}}\varepsilon_{n}^{1-\frac{2}{m}}.$ (B.6)
Therefore, observing that none of the upper bounds in (B.3)–(B.6) depend on $j$ and $k$ and that the one in (B.3) is the largest one, we bound each of the eight terms in (B.2) by this. Hence,
$\max_{j,k\in\mathinner{\{1,\ldots,d\}}}\mathinner{\!\bigl\lvert\overline{\Sigma}_{\varepsilon_{n},j,k}-\Sigma_{j,k}\bigr\rvert}$ $\leq 32\sigma_{m}^{2}\mathinner{\bigl(1+1/(1-\varepsilon_{n}-c^{-1}\varepsilon_{n})^{2/m}\bigr)}\mathinner{\Bigl(\frac{c+1}{c}\Bigr)}^{1-\frac{2}{m}}\varepsilon_{n}^{1-\frac{2}{m}}$
$\leq 32\sigma_{m}^{2}\mathinner{\Bigl(\frac{3c-1}{c-1}\Bigr)}\mathinner{\Bigl(\frac{c+1}{c}\Bigr)}\varepsilon_{n}^{1-\frac{2}{m}},$
the last estimate following from $\varepsilon_{n}\in(0,0.5)$$c>1$, and $m>2$. ∎
@@
Fix $c\in(1,\sqrt{1.5})$, and let Assumption 2.1 be satisfied with $m>2$. If $\varepsilon_{n}\in(0,0.5)$, with $\varepsilon_{n}$ as in (5), and there exists a strictly positive constants $\mathfrak{b}_{1}$ such that for all $j=1,\ldots,d$
$\mathbb{E}\mathinner{\bigl[\phi_{\underline{\alpha}_{j},\underline{\beta}_{j}}(X_{1,j})-\mathbb{E}\phi_{\underline{\alpha}_{j},\underline{\beta}_{j}}(X_{1,j})\bigr]}^{2}>\mathfrak{b}_{1}^{2}\qquad\text{and}\qquad\mathbb{E}\mathinner{\bigl[\phi_{\overline{\alpha}_{j},\overline{\beta}_{j}}(X_{1,j})-\mathbb{E}\phi_{\overline{\alpha}_{j},\overline{\beta}_{j}}(X_{1,j})\bigr]}^{2}>\mathfrak{b}_{1}^{2},$
@@
$C\mathinner{\biggl(\frac{12}{[(c-1)/c]^{1/m}}\biggr)}^{\frac{1}{2}}\sigma_{m}^{1/2}\mathinner{\biggl(\frac{\log^{5-\frac{2}{m}}(dn)}{n^{1-\frac{2}{m}}}\biggr)}^{\frac{1}{4}},$
where $C$ is a constant depending only on $\mathfrak{b}_{1}$ and $b_{2}$.
@@
We establish the first bound by verifying the conditions of Theorem 2.1 in Chernozhuokov et al. (2022), cf. in particular the display following that theorem. The second bound is proven analogously. First, note that for $j=1,\ldots,d$
$\phi_{\overline{\alpha}_{j},\overline{\beta}_{j}}(X_{i,j})-\mathbb{E}\phi_{\overline{\alpha}_{j},\overline{\beta}_{j}}(X_{i,j})\leq\overline{\beta}_{j}-\overline{\alpha}_{j}=Q_{1-\varepsilon_{n}+c^{-1}\varepsilon_{n}}(X_{1,j})-Q_{\varepsilon_{n}+c^{-1}\varepsilon_{n}}(X_{1,j}),$
Furthermore, by Lemma G.1 and $\varepsilon_{n}\geq\lambda_{2,c}(\delta,n)\frac{\log(dn)}{n}\geq\frac{\log(dn)}{3n}$, the far right-hand side of the previous display is upper bounded by
$\frac{\sigma_{m}}{(\varepsilon_{n}-c^{-1}\varepsilon_{n})^{1/m}}+\frac{\sigma_{m}}{(\varepsilon_{n}+c^{-1}\varepsilon_{n})^{1/m}}\leq\frac{2\sigma_{m}}{[(c-1)/c]^{1/m}\varepsilon_{n}^{1/m}}\leq\frac{6\sigma_{m}}{[(c-1)/c]^{1/m}}\mathinner{\biggl(\frac{n}{\log(dn)}\biggr)}^{1/m}.$
Similarly,
$\phi_{\overline{\alpha}_{j},\overline{\beta}_{j}}(X_{i,j})-\mathbb{E}\phi_{\overline{\alpha}_{j},\overline{\beta}_{j}}(X_{i,j})$ $\geq\overline{\alpha}_{j}-\overline{\beta}_{j}=Q_{\varepsilon_{n}+c^{-1}\varepsilon_{n}}(X_{1,j})-Q_{1-\varepsilon_{n}+c^{-1}\varepsilon_{n}}(X_{1,j})$
$\geq-\frac{6\sigma_{m}}{[(c-1)/c]^{1/m}}\mathinner{\biggl(\frac{n}{\log(dn)}\biggr)}^{1/m},$
@@
$B_{n}=\frac{12\sigma_{m}}{[(c-1)/c]^{1/m}}\mathinner{\biggl(\frac{n}{\log(dn)}\biggr)}^{1/m}.$
@@
a proof of the second inequality can be found in, e.g., Corollary 3 in Chow and Studden (1969). Therefore, by Theorem 2.1 and the display following it in Chernozhuokov et al. (2022), there exists a constant $C$ depending only on $\mathfrak{b}_{1}$ and $b_{2}$ such that
@@
$\leq$ $C\mathinner{\biggl(\frac{B_{n}^{2}\log^{5}(dn)}{n}\biggr)}^{\frac{1}{4}}=C\mathinner{\biggl(\frac{12}{[(c-1)/c]^{1/m}}\biggr)}^{\frac{1}{2}}\sigma_{m}^{1/2}\mathinner{\biggl(\frac{\log^{5-\frac{2}{m}}(dn)}{n^{1-\frac{2}{m}}}\biggr)}^{\frac{1}{4}}.$
@@
Fix $c\in(1,\sqrt{1.5})$, and let Assumption 2.1 be satisfied with $m>2$. If $\varepsilon_{n}\in(0,0.5)$, with $\varepsilon_{n}$ as in (5), then for $\overline{I}_{n,3}$ and $\underline{I}_{n,3}$ as defined in (A.6), it holds that
$\underline{I}_{n,3}\vee\overline{I}_{n,3}\leq\sigma_{m}f(c,m)\mathinner{\biggl(\overline{\eta}_{n}^{1-\frac{1}{m}}+\mathinner{\Bigl[\frac{\log(dn)}{n}\Bigr]}^{1-\frac{1}{m}}\biggr)},$
for a function $f\mathrel{\mathop{\ordinarycolon}}(1,\sqrt{1.5})\times[1,\infty)\to[0,\infty)$ of $c$ and $m$ only.
@@
Lemma G.4 and subadditivity of $z\mapsto z^{1-\frac{1}{m}}$ along with
$\varepsilon_{n}\leq\frac{c}{1-\sqrt{2(c^{2}-1)}}\cdot\overline{\eta}_{n}+\mathinner{\biggl[\frac{c}{3[1-\sqrt{2(c^{2}-1)}]}\vee c\mathinner{\biggl(\sqrt{2\frac{c+1}{c-1}}+\frac{1}{3}\biggr)}\biggr]}\frac{\log(dn)}{n}$
implies that for a non-negative function $f$ of $c$ and $m$ only
$\underline{I}_{n,3}\vee\overline{I}_{n,3}\leq\sigma_{m}f(c,m)\mathinner{\biggl(\overline{\eta}_{n}^{1-\frac{1}{m}}+\mathinner{\Bigl[\frac{\log(dn)}{n}\Bigr]}^{1-\frac{1}{m}}\biggr)}.$
The following lemma collects some simple limits calculations for later reference.
@@
$\log^{2}(d)\mathinner{\biggl(\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl[\frac{\log(dn)}{n}\Bigr]}^{\frac{1}{2}-\frac{1}{m}}\biggr)}\to 0,$ (B.9)
@@
$\sqrt{\log(d)\log(dn)}\mathinner{\biggl(\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl[\frac{\log(dn)}{n}\Bigr]}^{\frac{1}{2}-\frac{1}{m}}\biggr)}\to 0.$ (B.10)
@@
Next, (B.9) is true since we have already established $\log^{2}(d)\overline{\eta}_{n}^{1-\frac{2}{m}}\to 0$ and
$\log^{2}(d)\mathinner{\Bigl(\frac{\log(d)}{n}\Bigr)}^{\frac{1}{2}-\frac{1}{m}}=\frac{\mathinner{\bigl[\log(d)\bigr]}^{\frac{5}{2}-\frac{1}{m}}}{n^{\frac{1}{2}-\frac{1}{m}}}\to 0\iff\frac{\log(d)}{n^{\frac{m-2}{5m-2}}}\to 0,$
the latter convergence being true by assumption. Conclude, once more, by a subsequence argument.
Finally, we establish (B.10) by showing that any subsequence possesses a further subsequence along which (B.10) is true. Thus, fix a subsequence. In case there exists a subsequence thereof along which $n\leq d$, then it suffices to show that along this subsequence
$\log(d)\mathinner{\biggl(\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl[\frac{\log(d)}{n}\Bigr]}^{\frac{1}{2}-\frac{1}{m}}\biggr)}\to 0,$
@@
$\log(n)\mathinner{\biggl(\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl[\frac{\log(n)}{n}\Bigr]}^{\frac{1}{2}-\frac{1}{m}}\biggr)}\to 0.$
@@
$-S_{n,W,j}=n^{-1/2}\sum_{i=1}^{n}\mathinner{\bigl(\phi_{-\hat{\beta}_{j},-\hat{\alpha}_{j}}(-\tilde{X}_{i,j})-(-\mu_{j})\bigr)},\qquad j=1,\ldots,d.$
@@
and let $Z_{2}\sim~\mathsf{N}_{2d}(0,\Xi)=\mathrel{\mathop{\ordinarycolon}}\nu^{\prime}$. Hence, the version of Theorem 2.1 that replaces $\rho_{n,W}$ by $\rho_{n,W}^{os}$ (and which is assumed to hold true) applies to $(S_{n,W}^{\prime},-S_{n,W}^{\prime})^{\prime}$ and yields, for $C^{*}$ a constant depending only on $b_{1},b_{2},c$ and $m$, the upper bound
@@
By Lemma B.5 it suffices to prove (6) for hyperractangles $H$ on the form $H=\bigtimes_{j=1}^{d}[-\infty,t_{j}]$ with $t_{j}\in\overline{\mathbb{R}}$. To this end, by Lemma G.3,
$\underline{I}_{n,1}\vee\overline{I}_{n,1}\leq 2\mathinner{\biggl[\frac{1-\sqrt{2(c^{2}-1)}}{(c-1)}\biggr]}^{\frac{1}{m}}\sigma_{m}\overline{\eta}_{n}^{1-\frac{1}{m}}\mathrel{\mathop{\ordinarycolon}}=I_{n,1}$ (C.1)

### v4

@@ Appendix B Preparatory lemmas
Let $\varepsilon_{n}$ be as in (4) and recall the quantities defined in (A.3) and (A.4). Let $\overline{\Sigma}_{\varepsilon_{n}}$ and $\underline{\Sigma}_{\varepsilon_{n}}$ be the matrices with $(j,k)$th entry
@@
Let Assumption 2.1 be satisfied with $m>2$. If $\varepsilon_{n}$ is chosen as in (4) satisfying (A.5), then, for a positive constant $C$ depending only on $\lambda_{2}$ and $m$, it holds that
$\max_{j,k\in\mathinner{\{1,\ldots,d\}}}\mathinner{\Bigl(\mathinner{\!\bigl\lvert\overline{\Sigma}_{\varepsilon_{n},j,k}-\Sigma_{j,k}\bigr\rvert}\vee\mathinner{\!\bigl\lvert\underline{\Sigma}_{\varepsilon_{n},j,k}-\Sigma_{j,k}\bigr\rvert}\Bigr)}\leq C\sigma_{m}^{2}\varepsilon_{n}^{1-\frac{2}{m}}=\mathrel{\mathop{\ordinarycolon}}\mathfrak{T}_{n}^{(2)}.$ (B.1)
@@
$=\mathrel{\mathop{\ordinarycolon}}T_{1,j}+T_{2,j}+T_{3,j}.$
Thus, for $j,k\in\mathinner{\{1,\ldots,d\}}$, it follows from the Cauchy-Schwarz inequality that
@@
Recall that $\overline{\alpha}_{j}=Q_{c_{2,n}\varepsilon_{n}}(X_{1,j})$ such that $\mathbb{P}(X_{1,j}<\overline{\alpha}_{j})\leq c_{2,n}\varepsilon_{n}$. Thus, it follows by Lemma G.1, followed by Hölder’s inequality, that there exists a positive constant $C$ depending only on $\lambda_{2}$ and $m$ such that
$\mathbb{E}T_{2,j}^{2}$ $\leq\mathbb{E}(\mu_{j}+\sigma_{m}/(1-c_{2,n}\varepsilon_{n})^{1/m}-X_{1,j})^{2}\mathds{1}\mathinner{\bigl(X_{1,j}<\overline{\alpha}_{j}\bigr)}$
$\leq\mathinner{\Bigl(\mathbb{E}\mathinner{\!\bigl\lvert\mu_{j}-X_{1,j}+\sigma_{m}/(1-c_{2,n}\varepsilon_{n})^{1/m}\bigr\rvert}^{m}\Bigr)}^{2/m}\mathbb{P}(X_{1,j}<\overline{\alpha}_{j})^{1-\frac{2}{m}}$
$\leq\mathinner{\Bigl(2^{m}\sigma_{m}^{m}+2^{m}\sigma_{m}^{m}/(1-c_{2,n}\varepsilon_{n})\Bigr)}^{2/m}(c_{2,n}\varepsilon_{n})^{1-\frac{2}{m}}$
$\leq C\sigma_{m}^{2}\varepsilon_{n}^{1-\frac{2}{m}},$ (B.3)
the last inequality using that $c_{2,n}\varepsilon_{n}<1/2$ by Lemma G.2 and (A.5) and the same lemma also bounding $c_{2,n}$ from above in terms of $\lambda_{2}$ only.
By the same arguments, using $\mathbb{P}(X_{1,j}>\overline{\beta}_{j})=1-\mathbb{P}(X_{1,j}\leq\overline{\beta}_{j})\leq c_{1,n}\varepsilon_{n}$ and adjusting $C$, one also has that
$\mathbb{E}T_{3,j}^{2}\leq C\sigma_{m}^{2}\varepsilon_{n}^{1-\frac{2}{m}}.$ (B.4)
@@
Recallling that $\overline{\alpha}_{j}=Q_{c_{2,n}\varepsilon_{n}}(X_{1,j})$, it follows by Lemma G.1 and Hölder’s inequality, that
@@
$\leq\mathbb{E}\mathinner{\bigl(|X_{1,j}-\mu_{j}||\mu_{k}+\sigma_{m}/(1-c_{2,n}\varepsilon_{n})^{1/m}-X_{1,k}|\mathds{1}\mathinner{\bigl(X_{1,k}<\overline{\alpha}_{k}\bigr)}\bigr)}$
$\leq\mathbb{E}\mathinner{\bigl(|X_{1,j}-\mu_{j}||X_{1,k}-\mu_{k}|+|X_{1,j}-\mu_{j}|\sigma_{m}/(1-c_{2,n}\varepsilon_{n})^{1/m}\bigr)}\mathds{1}\mathinner{\bigl(X_{1,k}<\overline{\alpha}_{k}\bigr)}$
$\leq\mathinner{\Bigl(\mathbb{E}\mathinner{\bigl(|X_{1,j}-\mu_{j}||X_{1,k}-\mu_{k}|+|X_{1,j}-\mu_{j}|\sigma_{m}/(1-c_{2,n}\varepsilon_{n})^{1/m}\bigr)}^{\frac{m}{2}}\Bigr)}^{\frac{2}{m}}(c_{2,n}\varepsilon_{n})^{1-\frac{2}{m}}$
$\leq\mathinner{\Bigl(2^{m/2}(\sigma_{m}^{m}+\sigma_{m/2}^{m/2}\sigma_{m}^{m/2}/(1-c_{2,n}\varepsilon_{n})^{1/2}\Bigr)}^{2/m}(c_{2,n}\varepsilon_{n})^{1-\frac{2}{m}}$
$\leq C\sigma_{m}^{2}\varepsilon_{n}^{1-\frac{2}{m}},$ (B.5)
the last inequality using that $c_{2,n}\varepsilon_{n}<1/2$ by Lemma G.2 and (A.5) and the same lemma also bounding $c_{2,n}$ from above in terms of $\lambda_{2}$ only (adjust $C$ if needed).
@@
$\mathinner{\!\bigl\lvert\mathbb{E}T_{1,j}T_{3,k}\bigr\rvert}\leq C\sigma_{m}^{2}\varepsilon_{n}^{1-\frac{2}{m}}.$ (B.6)
Therefore, observing that none of the upper bounds in (B.3)–(B.6) depend on $j$ and $k$ it follows that (adjust $C$ if needed)
$\max_{j,k\in\mathinner{\{1,\ldots,d\}}}\mathinner{\!\bigl\lvert\overline{\Sigma}_{\varepsilon_{n},j,k}-\Sigma_{j,k}\bigr\rvert}\leq C\sigma_{m}^{2}\varepsilon_{n}^{1-\frac{2}{m}}.$
@@
Let Assumption 2.1 be satisfied with $m>2$. Let $\varepsilon_{n}$ be chosen as in (4) satisfying (A.5) and suppose that there exists a strictly positive constants $\mathfrak{b}_{1}$ such that for all $j=1,\ldots,d$
$\mathbb{E}\mathinner{\bigl[\phi_{\underline{\alpha}_{j},\underline{\beta}_{j}}(X_{1,j})-\mathbb{E}\phi_{\underline{\alpha}_{j},\underline{\beta}_{j}}(X_{1,j})\bigr]}^{2}>\mathfrak{b}_{1}^{2}\qquad\text{and}\qquad\mathbb{E}\mathinner{\bigl[\phi_{\overline{\alpha}_{j},\overline{\beta}_{j}}(X_{1,j})-\mathbb{E}\phi_{\overline{\alpha}_{j},\overline{\beta}_{j}}(X_{1,j})\bigr]}^{2}>\mathfrak{b}_{1}^{2}.$
@@
$C\sigma_{m}^{1/2}\mathinner{\biggl(\frac{\log^{5-\frac{2}{m}}(dn)}{n^{1-\frac{2}{m}}}\biggr)}^{\frac{1}{4}},$
where $C$ is a constant depending only on $\mathfrak{b}_{1}$$b_{2}$$\lambda_{1}$$\lambda_{2}$, and $m$.
@@
We establish the first bound by verifying the conditions of Theorem 2.1 in Chernozhukov et al. (2022), cf. in particular the display following that theorem. The second bound is proven analogously. First, note that for $j=1,\ldots,d$
$\phi_{\overline{\alpha}_{j},\overline{\beta}_{j}}(X_{i,j})-\mathbb{E}\phi_{\overline{\alpha}_{j},\overline{\beta}_{j}}(X_{i,j})\leq\overline{\beta}_{j}-\overline{\alpha}_{j}=Q_{1-c_{1,n}\varepsilon_{n}}(X_{1,j})-Q_{c_{2,n}\varepsilon_{n}}(X_{1,j}),$
Furthermore, by Lemma G.1 and $\varepsilon_{n}\geq\lambda_{2}\frac{\log(dn)}{n}$, the far right-hand side of the previous display is upper bounded by
$\frac{\sigma_{m}}{(c_{1,n}\varepsilon_{n})^{1/m}}+\frac{\sigma_{m}}{(c_{2,n}\varepsilon_{n})^{1/m}}\leq C_{1}\sigma_{m}\mathinner{\biggl(\frac{n}{\log(dn)}\biggr)}^{1/m},$
where, by Lemma G.2, $C_{1}$ is a constant depending only on $\lambda_{1}$$\lambda_{2}$ and $m$. Similarly,
$\phi_{\overline{\alpha}_{j},\overline{\beta}_{j}}(X_{i,j})-\mathbb{E}\phi_{\overline{\alpha}_{j},\overline{\beta}_{j}}(X_{i,j})$ $\geq\overline{\alpha}_{j}-\overline{\beta}_{j}=Q_{c_{2,n}\varepsilon_{n}}(X_{1,j})-Q_{1-c_{1,n}\varepsilon_{n}}(X_{1,j})$
$\geq-C_{1}\sigma_{m}\mathinner{\biggl(\frac{n}{\log(dn)}\biggr)}^{1/m},$
@@
$B_{n}=2C_{1}\sigma_{m}\mathinner{\biggl(\frac{n}{\log(dn)}\biggr)}^{1/m}.$
@@
a proof of the second inequality can be found in, e.g., Corollary 3 in Chow and Studden (1969). Therefore, by Theorem 2.1 and the display following it in Chernozhukov et al. (2022), there exists a constant $C_{2}$ depending only on $\mathfrak{b}_{1}$ and $b_{2}$ such that
@@
$\leq$ $C_{2}\mathinner{\biggl(\frac{B_{n}^{2}\log^{5}(dn)}{n}\biggr)}^{\frac{1}{4}}=C_{2}\cdot(2C_{1})^{1/2}\sigma_{m}^{1/2}\mathinner{\biggl(\frac{\log^{5-\frac{2}{m}}(dn)}{n^{1-\frac{2}{m}}}\biggr)}^{\frac{1}{4}},$
from which the conclusion of the lemma follows. ∎
@@
Let Assumption 2.1 be satisfied with $m>2$. If $\varepsilon_{n}$ is chosen as in (4) and satisfies (A.5), then for $\overline{I}_{n,3}$ and $\underline{I}_{n,3}$ as defined in (A.9), it holds that
$\underline{I}_{n,3}\vee\overline{I}_{n,3}\leq\sigma_{m}C\mathinner{\biggl(\overline{\eta}_{n}^{1-\frac{1}{m}}+\mathinner{\Bigl[\frac{\log(dn)}{n}\Bigr]}^{1-\frac{1}{m}}\biggr)},$
for a positive constant $C$ depending only on $\lambda_{1}$$\lambda_{2}$, and $m$.
@@
This is a trivial consequence of Lemma G.5 and subadditivity of $z\mapsto z^{1-\frac{1}{m}}$. ∎
The following lemma collects some simple limits for later reference.
@@
$\log^{2}(d)\mathinner{\biggl[\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl(\frac{\log(dn)}{n}\Bigr)}^{1-\frac{1}{(m/2)\wedge 2}}\biggr]}\to 0,$ (B.9)
@@
$\sqrt{\log(d)\log(dn)}\mathinner{\biggl(\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl[\frac{\log(dn)}{n}\Bigr]}^{1-\frac{1}{(m/2)\wedge 2}}\biggr)}\to 0.$ (B.10)
@@
Next to establish, (B.9) note that we have already shown that $\log^{2}(d)\overline{\eta}_{n}^{1-\frac{2}{m}}\to 0$. Furthermore, in case $m\leq 4$,
$\log^{2}(d)\mathinner{\Bigl(\frac{\log(d)}{n}\Bigr)}^{1-\frac{1}{(m/2)\wedge 2}}=\log^{2}(d)\mathinner{\Bigl(\frac{\log(d)}{n}\Bigr)}^{1-\frac{2}{m}}=\frac{\mathinner{\bigl[\log(d)\bigr]}^{3-\frac{2}{m}}}{n^{1-\frac{2}{m}}}\to 0\iff\frac{\log(d)}{n^{\frac{m-2}{3m-2}}}\to 0,$
the latter convergence being true because it is assumed that even $\log(d)/n^{\frac{m-2}{5m-2}}\to 0$. Conclude, once more, by a subsequence argument. The case of $m>4$ is handled analogously.
Finally, we establish (B.10) by showing that any subsequence possesses a further subsequence along which (B.10) is true. Thus, fix a subsequence. In case there exists a subsequence thereof along which $n\leq d$, it suffices to show that along this subsequence
$\log(d)\mathinner{\biggl(\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl[\frac{\log(d)}{n}\Bigr]}^{1-\frac{1}{(m/2)\wedge 2}}\biggr)}\to 0,$
@@
$\log(n)\mathinner{\biggl(\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl[\frac{\log(n)}{n}\Bigr]}^{1-\frac{1}{(m/2)\wedge 2}}\biggr)}\to 0.$
@@
$S_{n,W,j}=n^{-1/2}\sum_{i=1}^{n}\mathinner{\bigl(\phi_{-\hat{\beta}_{j},-\hat{\alpha}_{j}}(-\tilde{X}_{i,j})-(-\mu_{j})\bigr)},\qquad j=1,\ldots,d.$
@@
and let $Z_{2}\sim~\mathsf{N}_{2d}(0,\Xi)=\mathrel{\mathop{\ordinarycolon}}\nu^{\prime}$. Hence, the version of Theorem 2.1 that replaces $\rho_{n,W}$ by $\rho_{n,W}^{os}$ (and which is assumed to hold true) applies to $(S_{n,W}^{\prime},-S_{n,W}^{\prime})^{\prime}$ and yields, for $C^{*}$ a constant depending only on $b_{1},b_{2},\lambda_{1},\lambda_{2}$, and $m$, the upper bound
@@
By Lemma B.5 it suffices to prove (5) for hyperractangles $H$ of the form $H=\bigtimes_{j=1}^{d}[-\infty,t_{j}]$ with $t_{j}\in\overline{\mathbb{R}}$. Furthermore, recalling the definitions of $\mathfrak{T}_{n}^{(1)}$ and $\mathfrak{T}_{n}^{2}$ in (A.5) and (B.1), respectively, and using that $\log(dn)/n\leq\lambda_{2}^{-1}\varepsilon_{n}$, one has with $C$ being the constant from (B.1) that
$\mathfrak{T}_{n}^{(1)}+\mathfrak{T}_{n}^{(2)}\leq 2\varepsilon_{n}+\lambda_{2}^{-1}\varepsilon_{n}+\sqrt{\lambda_{2}^{-2}\varepsilon_{n}^{2}+4\lambda_{2}^{-1}\varepsilon_{n}^{2}}+Cb_{2}^{2}\varepsilon_{n}^{1-\frac{2}{m}}\leq C_{1}\varepsilon_{n}^{1-\frac{2}{m}},$ (C.1)
for $C_{1}$ a constant depending only on $b_{2}$$\lambda_{2}$ and $m$, and where we used that $\varepsilon_{n}\leq\varepsilon_{n}^{1-\frac{2}{m}}$ as $\varepsilon_{n}\in(0,0.5)$. In the following we first suppose that
$C_{1}\varepsilon_{n}^{1-\frac{2}{m}}<\frac{(b_{1}^{2}\wedge 1)}{2},$ (C.2)
which, by the penultimate display, implies that $\mathfrak{T}_{n}^{(1)}<1/2$ such that (A.5) is satisfied. By Lemma G.4, there exists a constant $C_{2}$ depending only on $\lambda_{1}$$\lambda_{2}$, and $m$ such that
$\underline{I}_{n,1}\vee\overline{I}_{n,1}\leq C_{2}\sigma_{m}\overline{\eta}_{n}^{1-\frac{1}{m}}\mathrel{\mathop{\ordinarycolon}}=I_{n,1}$ (C.3)